# Efficient Data-Driven Model Predictive Control for Demand Response of Commercial Buildings

### v2

@@ I Introduction
The built environment stands as one of the most energy-intensive sectors, and as one of the main contributors to climate change. The buildings sector alone accounts for 30% of global energy consumption and contributes to 28% of total greenhouse gas emissions [1]. In the U.S., commercial buildings (CBs) alone consume 46% of the energy used in the building sector. A substantial portion of this energy is allocated to operating their heating, ventilation, and air conditioning (HVAC) systems, which constitute 47% of their total energy consumption [2]. Inefficient HVAC operation, due to the lack of adequate control, often results in significant energy waste [3]. In the U.S., over 90% of CBs fall into the small- to medium-sized category and are predominantly equipped with packaged rooftop units (RTUs). It is estimated that the majority of these buildings still rely on rudimentary controllers, such as ruled-based or set-point (SP)-based [4]. Model predictive control (MPC) has demonstrated its ability to improve thermal comfort while achieving energy and cost savings through efficient HVAC operation [3]. Despite extensive research and practical case studies, adoption of the technology within the CBs industry remains limited, primarily due to challenges in automating design, tuning, and implementation. The quality of an MPC solution depends on the model’s accuracy and the ability to solve the optimization problem efficiently and globally at its core. Physics-based approaches produce reliable, interpretable models but require significant development effort [5]. In contrast, data-driven approaches promote wide-scale implementation, because they do not require understandings of the system dynamics. This translates into a more transferable approach and lower development costs [3]. Modelling the thermal dynamics of a building can be a challenging task due to their highly nonlinear nature [6]. To address these challenges, many have adopted an approach based on recurrent neural networks (RNNs) because of their high modelling accuracy for dynamic systems [6]. Although RNN-based MPC have been applied to CBs in [7], this results in a nonconvex optimization problem which is computationally intractable. In practice, modelling accuracy is often sacrificed for control tractability, favoring linear models despite their limitations [6].
@@
    We leverage a precise discrete model to formulate an adaptable real-time DR bidding strategy (Section III-B2).
@@ II-C1 Base Controller
$\sum_{\tau=k-\rho}^{k-1}\max\left\{\mathbf{u}^{\tau-1}-\mathbf{u}^{\tau}\right\}\leq 1-\mathbf{u}^{k},\quad$ $\forall k\in\mathcal{T}_{t}$ (6)
@@
We consider a time-varying rates for energy $\lambda_{\text{E}}^{t}\geq 0$, and a fixed rate on peak power demand $\lambda_{\text{P}}\geq 0$, as utilities usually bill commercial consumers for both. Peak demand fees are based on the maximum power usage observed during a month [18]. Because of the lack of whole building’s power profiles due, for example, to submetering, we aim to minimize the peak demand induced by the RTUs over the prediction horizon as a proxy. Constraints (4)--(5) represent the thermal comfort zone, where $\bar{\boldsymbol{\theta}}^{t}\in\mathbb{R}^{n_{\text{z}}}$, $t\in\{1,2,\dots\}$, is the upper bound on IAT, for all $n_{\text{z}}$ zones. The building’s thermal dynamics are described by a parametrized ICRNN, $f_{\text{ICRNN}}:\mathbb{R}^{(2n_{\text{u}}n_{\text{z}}+n_{\text{s}})\times T}\mapsto\mathbb{R}^{n_{\text{z}}\times T}$ that takes as inputs extended controls $\hat{\mathbf{u}}^{t}\in\mathcal{\hat{U}}$, and the $n_{s}\in\mathbb{N}$ building’s state variables, i.e., exogenous variables, defined for all $t\in\{1,2,\dots\}$, as $\mathbf{s}^{t}\in\mathbb{R}^{n_{\text{s}}}$. Details about the ICRNN, including its architecture and training procedure, are presented in Section III-A1. We omit the lower bound on IAT $\underaccent{\bar}{\boldsymbol{\theta}}^{t}\in\mathbb{R}^{n_{\text{z}}}$, $t\in\{1,2,\dots\}$, to maintain convexity as superlevel sets of a convex functions are generally non-convex [12]. This relaxation is not limiting for cooling operations because, by minimizing energy consumption and peak power, the IAT will tend to be close to the upper bound. We plan to extend this approach to heating operations in future work. Constraint (6) prevents equipment toggling, as short cycling significantly decreases the efficiency of units [19]. It ensures that each RTU component remains in the same state during $\rho\in\mathbb{N}$ time steps when transitioning from on to off. Here, $\mathbf{0}\in\mathbb{R}^{n_{\text{z}}n_{\text{u}}}$ denotes the zero vector. Constraint (7) represents the energy consumed during a control round. Constraint (12) defines the inputs of the ICRNN, and (13) states the domains of the optimization variables. Considering our ICRNN architecture further described in Section III-A1, the problem (3)--(13) is a mixed-integer piecewise linear optimization problem. When the binary variables are relaxed to continuous ones, its objective function and its feasible set are convex, and the proposed MPC is a convex problem [20]. Convexity of the relaxed optimization problem guarantees that the solver avoids getting stuck in local minima, and therefore promotes higher solution quality with respect to conventional NN-based MPC.
@@ III-A2 Benchmark
We use a multi-output first order autoregressive exogenous input (ARX) model and employ a recursive approach for generating multi-step-ahead prediction. Input features include OAT, time of day of week in sin/cos encoding at $t$, GHI from $t-3$ to $t$, and $\Delta\boldsymbol{\theta}^{t}$ from $t-4$ to $t-1$.
@@ III-A3 Results
We compare our modelling approach with benchmark models. Figure 2 illustrates the prediction performance over a 24-hour period in one of the two zones. To fit the NN models, we employ a recurrent layer of 60 units for the ICRNN and 45 units for the LSTM. We compute the corresponding root mean square error (RMSE) and their standard deviation for each predictive model. In terms of fitting performances, the linear model produces the largest error values with a RMSE of ($0.45\pm 0.04$)C. This is primarily due to error propagation over time resulting from the recursive approach and its limitation in capturing nonlinear patterns, such as coupling between zones and thermal inertia. The LSTM model, conversely, shows the lowest error values with a RMSE of ($0.09\pm 0.01$)C. An LSTM NN has a memory cell that can capture information from earlier time steps and store it for later use, which helps addressing the problem of vanishing or exploding gradient [17], thereby contributing to a more accurate representation of thermal inertia. Our ICRNN model shows significant improvements when compared to the linear model with a RMSE of ($0.14\pm 0.03$)C. Because the ICRNN is a piecewise linear convex approximation, it has representation limitations. It remains a competitive alternative to the LSTM.
@@ IV Conclusion
In this work, we propose an energy management method for RTU-HVAC of small CBs. Our convex MPC approach levrages a discrete and data-driven model of the thermal dynamics, making the implementation in RTU-based CBs straightforward. We develop a bidding strategy for the real-time market and apply our approach to a TOU and a CPR program. The performance of the proposed controller is evaluated through a numerical building simulation, and has demonstrated a good compromise between thermal comfort, energy reduction, and cost. In future work, we will extend our approach to an aggregation of small CBs or to a large CB through a decentralized optimization method. We also plan to propose a resolution method that produces provably optimal solutions.

### v3

@@ I Introduction
The built environment stands as one of the most energy-intensive sectors, and as one of the main contributors to climate change. The buildings sector alone accounts for 30% of global energy consumption and contributes to 28% of total greenhouse gas emissions [1]. In the U.S., commercial buildings (CBs) alone consume 46% of the energy used in the building sector. A substantial portion of this energy is allocated to operating their heating, ventilation, and air conditioning (HVAC) systems, which constitute 47% of their total energy consumption [2]. Inefficient HVAC operation, due to the lack of adequate control, often results in significant energy waste [3]. In the U.S., over 90% of CBs fall into the small- to medium-sized category and are predominantly equipped with packaged rooftop units (RTUs). It is estimated that the majority of these buildings still rely on rudimentary controllers, such as ruled-based or set-point (SP)-based.
Model predictive control (MPC) has demonstrated its ability to improve thermal comfort while achieving energy and cost savings through efficient HVAC operation [3]. Despite extensive research and practical case studies, adoption of the technology within the CBs industry remains limited, primarily due to challenges in automating design, tuning, and implementation. The quality of an MPC solution depends on the model’s accuracy and the ability to solve the optimization problem efficiently and globally at its core. Physics-based approaches produce reliable, interpretable models but require significant development effort [4]. In contrast, data-driven approaches promote wide-scale implementation, because they do not require understandings of the system dynamics. This translates into a more transferable approach and lower development costs [3]. Modelling the thermal dynamics of a building can be a challenging task due to their highly nonlinear nature [5]. To address these challenges, many have adopted an approach based on recurrent neural networks (RNNs) because of their high modelling accuracy for dynamic systems [6]. Although RNN-based MPC have been applied to CBs in [7], this results in a nonconvex optimization problem which is computationally intractable. In practice, modelling accuracy is often sacrificed for control tractability, favoring linear models despite their limitations [6].
@@
    We leverage a precise discrete model to formulate an adaptable real-time DR bidding strategy (Section II-C2).
@@ II-C1 Base Controller
$\sum_{\tau=k-\rho}^{k-1}\max\left\{\mathbf{u}^{\tau-1}-\mathbf{u}^{\tau},\mathbf{0}\right\}\leq 1-\mathbf{u}^{k},\forall k\in\mathcal{T}_{t}$ (6)
@@
We consider a time-varying rates for energy $\lambda_{\text{E}}^{t}\geq 0$, and a fixed rate on peak power demand $\lambda_{\text{P}}\geq 0$, as utilities usually bill commercial consumers for both. Peak demand fees are based on the maximum power usage observed during a month [19]. Because of the lack of whole building’s power profiles due, for example, to submetering, we aim to minimize the peak demand induced by the RTUs over the prediction horizon as a proxy. Constraints (4)--(5) represent the thermal comfort zone, where $\bar{\boldsymbol{\theta}}^{t}\in\mathbb{R}^{n_{\text{z}}}$, $t\in\{1,2,\dots\}$, is the upper bound on IAT, for all $n_{\text{z}}$ zones. The building’s thermal dynamics are described by a parametrized ICRNN, $f_{\text{ICRNN}}:\mathbb{R}^{(2n_{\text{u}}n_{\text{z}}+n_{\text{s}})\times T}\mapsto\mathbb{R}^{n_{\text{z}}\times T}$ that takes as inputs extended controls $\hat{\mathbf{u}}^{t}\in\mathcal{\hat{U}}$, and the $n_{s}\in\mathbb{N}$ building’s state variables, i.e., exogenous variables, defined for all $t\in\{1,2,\dots\}$, as $\mathbf{s}^{t}\in\mathbb{R}^{n_{\text{s}}}$. Details about the ICRNN, including its architecture and training procedure, are presented in Section III-A1. We omit the lower bound on IAT $\underaccent{\bar}{\boldsymbol{\theta}}^{t}\in\mathbb{R}^{n_{\text{z}}}$, $t\in\{1,2,\dots\}$, to maintain convexity as superlevel sets of (4) are not garanteed to be convex [12]. This relaxation is not limiting for cooling operations because, by minimizing energy consumption and peak power, the IAT will tend to be close to the upper bound. We plan to extend this approach to heating operations in future work. Constraint (6) prevents equipment toggling, as short cycling significantly decreases the efficiency of units [20]. It ensures that each RTU component remains in the same state during $\rho\in\mathbb{N}$ time steps when transitioning from on to off. Here, $\mathbf{0}\in\mathbb{R}^{n_{\text{z}}n_{\text{u}}}$ denotes the zero vector. Constraint (7) represents the energy consumed during a control round. Constraint (12) defines the inputs of the ICRNN, and (13) states the domains of the optimization variables. Considering our ICRNN architecture further described in Section III-A1, the problem (3)--(13) is a mixed-integer piecewise linear optimization problem. When the binary variables are relaxed to continuous ones, its objective function and its feasible set are convex, and the proposed MPC is a convex problem [21]. Convexity of the relaxed optimization problem guarantees that the solver avoids getting stuck in local minima, and therefore promotes higher solution quality with respect to conventional NN-based MPC.
@@ III-A2 Benchmark
We use a multi-output first-order autoregressive exogenous input (ARX) model similar to [27] and employ a recursive approach for generating multi-step-ahead prediction. Input features include OAT, time of day of week in sin/cos encoding at $t$, GHI from $t-3$ to $t$, and $\Delta\boldsymbol{\theta}^{t}$ from $t-4$ to $t-1$.
@@ III-A3 Results
We compare our modelling approach with benchmark models. Figure 2 illustrates the prediction performance over a 24-hour period in one of the two zones. To fit the NN models, we employ a recurrent layer of 60 units for the ICRNN and 45 units for the LSTM. We compute the corresponding root mean square error (RMSE) and their standard deviation for each predictive model. In terms of fitting performances, the linear model produces the largest error values with a RMSE of ($0.45\pm 0.04$)C. This is primarily due to error propagation over time resulting from the recursive approach and its limitations in capturing nonlinear patterns such as coupling between zones [5]. The LSTM model, conversely, shows the lowest error values with a RMSE of ($0.09\pm 0.01$)C. An LSTM NN has a memory cell that can capture information from earlier time steps and store it for later use, which helps addressing the problem of vanishing or exploding gradient [18], thereby contributing to a more accurate representation of thermal inertia. Our ICRNN model shows significant improvements when compared to the linear model with a RMSE of ($0.14\pm 0.03$)C. Because the ICRNN is a piecewise linear convex approximation, it has representation limitations. It remains a competitive alternative to the LSTM.
@@ IV Conclusion
In this work, we propose an energy management method for RTU-HVAC of small CBs. Our convex MPC approach levrages a discrete and data-driven model of the thermal dynamics, making the implementation in RTU-based CBs straightforward. We develop a bidding strategy for the real-time market and apply our approach to a TOU and a CPR program. The performance of the proposed controller is evaluated through a numerical building simulation, and has demonstrated a good compromise between thermal comfort, energy reduction, and cost. In future work, we will extend our approach to an aggregation of small CBs or to a large CB through distributed optimization via the alternating direction method of multipliers (ADMM) [39]. We also plan to propose a resolution method that produces provably optimal solutions.